OpenStudy (anonymous):

What is the area of a triangle that has a height of 10 feet and a base length of 8 feet? 10 feet2 20 feet2 40 feet2 80 feet2

OpenStudy (michele_laino):

here we have to apply the subsequent formula: \[area = \frac{{base \times height}}{2} = \frac{{10 \times 8}}{2} = ...?\]

OpenStudy (anonymous):

alright let me write that down (:

OpenStudy (anonymous):

Ok so I got 40

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

so C?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Alright, I have one more

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

Zeda described four triangles as shown below: Triangle P: Two angles measure 50°. Triangle Q: All sides have length 13 cm. Triangle R: Two sides have length 3 cm, and the included angle measures 60°. Triangle S: Base has length 5 cm, and base angles measure 30°. Which triangle is not a unique triangle? Triangle P Triangle Q Triangle R Triangle S

OpenStudy (michele_laino):

please look at this drawing: 

OpenStudy (michele_laino):

Is my drawing possible?

OpenStudy (anonymous):

I think so.. yes

OpenStudy (michele_laino):

yes! In fact i can draw infinity triangles which have two angles of 50 degrees. So, what is the right option?

OpenStudy (anonymous):

Triangle P?

OpenStudy (michele_laino):

I meant an infinite number of such triangles. Correct!

OpenStudy (anonymous):

Okay Thank you so Much!

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

Do you think you can help with one more, I didnt see this question.. I promise last one lol @Michele_Laino

OpenStudy (anonymous):

@Nnesha